OpenStudy (anonymous):

For the simple harmonic motion equation d=2sin((π/3)t), what is the frequency?

OpenStudy (anonymous):

π/3=ω, ω=2πf

OpenStudy (anonymous):

π/3=2πf, π/6=πf, f=1/6